OpenStudy (gracygirl):

Write the point-slope form of the equation of the line that passes through the origin and has a slope of 2. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution. @Will.H

OpenStudy (gracygirl):

@Awolflover1 @AloneS

OpenStudy (will.h):

So we have a slope of 2 and the line passes through (0,0) As we solved before we have to use the following formula y-y1 = m(x - x1) We know that y1 is 0 and so is x1 So substitute y - 0 = 2(x - 0) That's in the point slope form Hope that helps